OpenStudy (anonymous):

Select the equation of the line that passes through the point (2, –1) and has slope 3 in point-slope form. A. (y – 1) = 3(x + 2) B. (y + 1) = 3(x – 2) C. (x + 2) = 3(y – 1) D. (x – 2) = 3(y + 1)

OpenStudy (anonymous):

y = mx + b Plug in (2, -1) and m = 3 -1 = 3(2) + b -7 = b y = 3x - 7 So looking at the multiple choices: A. (y – 1) = 3(x + 2) is the same as y = 3x + 7 B. (y + 1) = 3(x – 2) is the same as y = 3x - 7 So B is the answer.

OpenStudy (anonymous):

thank you!